OpenStudy (anonymous):

Which shows the prime factorization of 36? A. 2 • 2 • 2 • 3 B. 2 • 2 • 3 • 3 C. 2 • 2 • 9 • 1 D. 2 • 3 • 3 • 3

OpenStudy (anonymous):

@brookie5150

OpenStudy (anonymous):

@Best_Mathematician

OpenStudy (anonymous):

B. 2.2.3.3 as 2*2=4*3=12*3=36